THE MILLIMETER/SUBMILLIMETER SPECTRUM OF LISH(X̃^{1}A′): FURTHER INVESTIGATIONS OF THE METAL-SULFUR BOND

\underline{A. JANCZYK} and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ 85721.

The pure rotational spectra of LiSH (X̃^{1}A′) and two of its isotopomers (^{6}LiSH, LiSD) were recorded using millimeter/submillimeter-wave direct absorption techniques. The molecule was produced by reacting lithium vapor with H_{2}S in the presence of a d.c. discharge. Transition ranging from J = 1 → 2 to J = 13 → 14, spanning the region of 73-520 GHz, were measured for ^{7}LiSH, as well as 3-4 transitions for the ^{6}Li and deuterium isotopomers. The spectra showed an extensive K_{a} ladder structure for the three species, which is consistent with the molecule being an asymmetric top. Rotational constants for all three molecules have been determined, as well as r_{0} and r_{m}^{(1)} structures for LiSH. This study shows that this molecule is bent, with Θ = 93^{∘} and therefore is significantly different from the quasilinear LiOH. This structural change likely results from more covalent bonding in LiSH.